# Design, Construction, and Test of Compact, Distributed-Charge, X-Band Accelerator Systems that Enable Image-Guided, VHEE FLASH Radiotherapy

### v1

@@ 8 A Vision for Image-Guided VHEE Radiation Therapy
The concomitant, inherently colinear production of high-brightness x-rays through laser-Compton scattering and VHEEs provides a unique opportunity for in-line image-guided radiation therapy using a single device. The notion that a Compton-scattered x-ray beam could be used for image guidance for an underlying VHEE source was, to the best of our knowledge, first suggested by DesRosiers, et al. [74] in the context of laser-plasma VHEE sources. Although DesRosiers, et al. (2008) [74] pointed to the potential benefits in x-ray quality, comparing laser-Compton x-rays to those produced by synchrotron facilities, there have not been, to the best of our knowledge, any discussions on the feasibility of a laser-Compton x-ray image-guided VHEE radiation therapy system. In this section, we aim to outline the challenges and propose solutions for implementing such a system.
Fundamentally, this concept requires the rapid switching between the production of laser-Compton x-rays to the delivery of a VHEE beam. The most straightforward implementation of this switching, considering the DCCS architecture, would be to shutter both ILS and PGL lasers, turn off the final electron bending magnet, and then release bursts of PGL pulses based on the prescribed radiation treatment. In this operation mode, the minimum time between imaging and therapy is the time required to operate laser shutters and the decay time of the bending magnet, which may be on the order of seconds. For this approach to work, the electron energy used to produce LCS x-rays must also be of an energy relevant for the radiation treatment. For the DCCS architecture presented here, producing a 70 keV imaging beam, which has been previously identified as an energy of interest for minimizing dose during phase contrast imaging applications [56], requires 45 MeV electrons. Recent treatment planning simulation studies have suggested that this energy range may be of interest, especially when operated at FLASH-relevant dose rates, for treating pediatric brain tumors [75]. However, 45 MeV electrons are insufficient for use in deep-seated tumors in adult humans, and simply increasing the electron energy will correspond with an x-ray beam whose energy is too high for practical clinical imaging applications. One of the appeals of producing x-rays or -rays through laser-Compton scattering is to minimize the energy requirements on the electron beam, thus enabling the use of compact accelerator architectures, especially when compared to the required synchrotron facility requirements to produce the same x-ray or -ray energy. Additionally, counter-propagating electron and laser beams uses the least energetic electrons possible for a given scattered x-ray energy.
@@
Another consideration for image-guided VHEE FLASH radiation therapy is identifying appropriate combinations of x-ray and electron beam transverse sizes for imaging and therapy respectively. Previously presented work has investigated the feasibility of using the free propagating x-ray and electron beams after passing through a common vacuum exit window, and has identified operating conditions of 37 MeV electrons and 51 keV Compton-scattered x-rays to potentially image and subsequently irradiate a target the size of a mouse skull [76]. 51 keV was chosen as the imaging energy to utilize a Gd foil to produce a K-edge hole that can be used as an image-guidance crosshair to more accurately identify the central propagation axis of the electron beam (see Figure 10). To fully explore the parameter space available, the incorporation of electron beam optics, x-ray optics, and considerations of different laser-electron interaction geometries should be considered, and will be explored further in future work.
@@ 9 Conclusion
Very high energy electrons (VHEEs) have been identified as a promising ionizing radiation modality, especially in the context of eliciting the FLASH effect, but VHEE source size and access to high-fidelity image guidance limit clinical implementation. In this work, we described the Distributed Charge Compton Source (DCCS) as a uniquely suited architecture for image-guided VHEE FLASH radiation therapy. Through maximally distributing electron charge in long, low-charge trains of microbunches at 11.424 GHz (X-band), large currents can be accelerated without compromising the quality of the electron beam. This conservation of electron beam quality enables the electron beam to be used not only directly as an ionizing radiation source, but also as a generator of secondary x-rays through the head-on collision with a counter-propagating train of laser pulses through the process of laser-Compton scattering. Since the laser micropulses and electron microbunches are generated using the same RF master lock, the pulse/bunch trains are inherently synchronized up to a simple phase delay. As a systems integration test, the DCCS has demonstrated the production of laser-Compton x-rays with narrow energy bandwidths (54 keV with 0.41% on-axis RMS bandwidth presented here) and the production of electron macrobunches with FLASH-relevant charge densities (14 nC in 86.6 ns at 100 Hz). As the underlying systems are continued to be commissioned, a primary milestone is to use the same electron beam for both the production of inherently colinear laser-Compton x-rays and electron irradiation in a single experimental session, thus demonstrating the feasibility of using the DCCS architecture for x-ray image-guided VHEE FLASH radiation therapy.
@@
CPJB, JMA, AJA, JCRB, SMB, MMC, RADLL, DAD, HHE, RF, AG, KJG, ASG, FVH, YH, GI, MJ, CAJ, KWK, AL, RJL, MWM, EM, CLN, HJN, KRP, ZRP, MEQ, FR, TR, JS, MES, MWLS, JYY, CBZ, LEZ, and EJZ declare that they have stock-related financial interest in Lumitron Technologies, Inc. CPJB holds patents related to the underlying research presented in this work.

### v2

@@ 8 A Vision for Image-Guided VHEE Radiation Therapy
The concomitant, inherently colinear production of high-brightness x rays through laser-Compton scattering and VHEEs provides a unique opportunity for in-line image-guided radiation therapy using a single device. The notion that a Compton-scattered x-ray beam could be used for image guidance for an underlying VHEE source was, to the best of our knowledge, first suggested by DesRosiers, et al. [84] in the context of laser-plasma VHEE sources. Although DesRosiers, et al. (2008) [84] pointed to the potential benefits in x-ray quality, comparing laser-Compton x rays to those produced by synchrotron facilities, there have not been, to the best of our knowledge, any discussions on the feasibility of a laser-Compton x-ray image-guided VHEE radiation therapy system. In this section, we aim to outline the challenges and propose solutions for implementing such a system.
Fundamentally, this concept requires the rapid switching between the production of laser-Compton x rays to the delivery of a VHEE beam. The most straightforward implementation of this switching, considering the DCCS architecture, would be to shutter both ILS and PGL lasers, turn off the final electron bending magnet, and then release bursts of PGL pulses based on the prescribed radiation treatment. In this operation mode, the minimum time between imaging and therapy is the time required to operate laser shutters and the decay time of the bending magnet, which may be on the order of seconds. For this approach to work, the electron energy used to produce LCS x rays must also be of an energy relevant for the radiation treatment. For the DCCS architecture presented here, producing a 70 keV imaging beam, which has been previously identified as an energy of interest for minimizing dose during phase contrast imaging applications [66], requires 45 MeV electrons. Recent treatment planning simulation studies have suggested that this energy range may be of interest, especially when operated at FLASH-relevant dose rates, for treating pediatric brain tumors [85]. However, 45 MeV electrons are insufficient for use in deep-seated tumors in adult humans, and simply increasing the electron energy will correspond with an x-ray beam whose energy is too high for practical clinical imaging applications. One of the appeals of producing x rays or rays through laser-Compton scattering is to minimize the energy requirements on the electron beam, thus enabling the use of compact accelerator architectures, especially when compared to the required synchrotron facility requirements to produce the same x-ray or -ray energy. Additionally, counter-propagating electron and laser beams uses the least energetic electrons possible for a given scattered x-ray energy.
@@
Another consideration for image-guided VHEE FLASH radiation therapy is identifying appropriate combinations of x-ray and electron beam transverse sizes for imaging and therapy respectively. For example, the 38-MeV electron beam considered in the manuscript in Section 6 resulted in a usable x-ray imaging field of about 8 cm at a distance of 5.66 m from the laser-electron interaction point. Although this field-of-view is smaller than what is used with current clinical x-ray sources, a benefit of using a low-divergence beam is that the detector can be placed further away from the patient while preserving the x-ray image. Scattered x rays from the patient are naturally filtered away, improving the image quality on the detector. Additionally, placing the detector further away from the patient enables phase-based image techniques, which can improve the differentiation of soft tissue for increased diagnostic potential. For the electron beam, preclinical small-animal studies are a primary short-term objective to demonstrate the feasibility of UHDR electron beam irradiation using the DCCS. Beams of this type can be constructed using natural beam expansion from the exit window and simple collimators and scattering foils [86]. For clinical applications, VHEE pencil-beam scanning is a potential modality, and a proposed magnet kicker system has recently been discussed [48]. However, pencil-beam scanning would result in an “image-informed” as opposed to a strictly “image-guided” modality in which the laser-Compton x-ray image can define the extent of the scan, but is no longer inherently colinear with the electron beam since the electron beam path is modified by a steering magnet. In general, as long as any electron beam shaping does not modify the central trajectory of the electron beam, the potential for x-ray image guidance is preserved.
Previously presented work has investigated the feasibility of using the free propagating x-ray and electron beams after passing through a common vacuum exit window, and has identified operating conditions of 37 MeV electrons and 51 keV Compton-scattered x rays to potentially image and subsequently irradiate a target the size of a mouse skull [87]. 51 keV was chosen as the imaging energy to utilize a Gd foil to produce a K-edge hole that can be used as an image-guidance crosshair to more accurately identify the central propagation axis of the electron beam (see Figure 10). To fully explore the parameter space available, the incorporation of electron beam optics, x-ray optics, and considerations of different laser-electron interaction geometries should be considered, and will be explored further in future work.
@@ 9 Conclusion
Very high energy electrons (VHEEs) have been identified as a promising ionizing radiation modality, especially in the context of eliciting the FLASH effect, but VHEE source size and access to high-fidelity image guidance limit clinical implementation. In this work, we described the Distributed Charge Compton Source (DCCS) as a uniquely suited architecture for image-guided VHEE FLASH radiation therapy. Through maximally distributing electron charge in long, low-charge trains of microbunches at 11.424 GHz (X-band), large currents can be accelerated without compromising the quality of the electron beam. This conservation of electron beam quality enables the electron beam to be used not only directly as an ionizing radiation source, but also as a generator of secondary x rays through the head-on collision with a counter-propagating train of laser pulses through the process of laser-Compton scattering. Since the laser micropulses and electron microbunches are generated using the same RF master lock, the pulse/bunch trains are inherently synchronized up to a simple phase delay. As a systems integration test, the DCCS has demonstrated the production of laser-Compton x rays with narrow energy bandwidths (54 keV with 0.41% on-axis RMS bandwidth presented here) and the production of electron macrobunches with FLASH-relevant charge densities (14 nC in 86.6 ns at 100 Hz). As the underlying systems are continued to be commissioned, a primary milestone is to use the same electron beam for both the production of inherently colinear laser-Compton x rays and electron irradiation in a single experimental session, thus demonstrating the feasibility of using the DCCS architecture for x-ray image-guided VHEE FLASH radiation therapy.
Ethics statement
Ethical approval was not required for the study involving animals in accordance with the local legislation and institutional requirements because the animal product (dried anchovy) used for imaging tests was purchased at a grocery store.
@@
CPJB, JMA, AJA, JCRB, SMB, MMC, RADLL, DAD, HHE, RF, AG, KJG, ASG, FVH, YH, GI, MJ, CAJ, KWK, AL, RJL, MWM, EM, CLN, HJN, KRP, ZRP, MEQ, FR, TR, JS, MES, MWLS, JYY, CBZ, LEZ, and EJZ were employed by Lumitron Technologies, Inc. CPJB holds patents related to the underlying research presented in this work.